OpenStudy (anonymous):

What's the perimiter of a square when its sides are 3x+6 and 5x?

OpenStudy (anonymous):

2(3x+6) + 2(5x) = 16x +12

OpenStudy (saifoo.khan):

Perimeter = 2(L+B) = 2( (3x+6) + (5x)) = 2( 8x+6) = 16x+12

OpenStudy (anonymous):

but since its a square, x = 3 so the perimeter is 60 :)

OpenStudy (anonymous):

why is x=3? bcuz its a square --> 3x +6 = 5x --> x=3

OpenStudy (anonymous):

So x=3?

OpenStudy (anonymous):

ya....and the perimeter is 60